# Efficient simulation of the SABR model

### v1

@@ 5 Numerical Results
In our numerical tests, unless otherwise specified, we used $N=10^{5}$ paths for each simulation, and repeated $m=50$ times. Out of $m$ prices, we computed the average bias (used the FDM price as the benchmark) and standard deviation. Our algorithm is implemented in MATLAB R2023b on a laptop with a 13th Gen Intel Core i7-1360P 2.20 GHz processor. In Section 5.2, to demonstrate the speed of algorithm, we compare the CPU time of our method with those reported in Cai et al. (2017). Cai et al. (2017) used Matlab 7 on an Intel Core2 Q9400 2.66GHZ processor. One may argue that comparing CPU time in this manner may not be fair as we used a newer system. However, the two CPUs are not significantly different to overturn the conclusion. As shown, our algorithm is faster than the existing algorithms by one or two orders of magnitude.
5.1 Accuracy and comparison to analytic approximations
In this subsection, we test the accuracy of our algorithm (Algorithm 2) using Cases I and II in Table 1. Tables 2 and 3 show the numerical results for the two sets. The two examples are previously tested in Antonov and Spector (2012) to demonstrate the performance of the cutting-edge analytic approximations for European vanilla options, map to the zero correlation (ZC Map, and hybrid map to the zero correlation (Hyb ZC Map), against the widely used Hagan et al. (2002)’s implied volatility formula. [^4] Since these approximation formulas are all expressed in terms of implied volatilities, we use the Black-Scholes formula to obtain the call option price. Antonov and Spector (2012) use MC method for the benchmark option price. Instead, we use FDM price for the benchmark, which is more precise. Therefore, we also display those option prices for comparison.
Tables 2 and 3 demonstrate that our simulation scheme is highly accurate. As we decrease the time step $h$ from 1 to $1/16$, the option price converges to its true price. Even for a large time step ($h=1$), the option prices are accurate almost up to three decimal points. Such bias level is comparable to the advanced analytic methods. Overall speaking, we note that the bias in Table 3 (Case II) is lower than that in Table 2 (Case I). This is probably due to the observation that the geometric BM approximation for the conditional mean $\bar{F}_{0}^{T}$ in Eq. (12) holds better as $\beta=0.6$ (Case II) is closer to one than $\beta=0.3$ (Case I).
@@ 5.1 Accuracy and comparison to analytic approximations
Bias of our simulation method  (×$(\times 10^{-3})$
$h=1$1.22 -1.49 -0.37 0.49 1.28 1.72 1.32
$h=1/4$0.46 -0.24 0.22 0.42 0.56 0.56 0.48
$h=1/16$0.34 -0.20 0.00 0.05 0.11 0.10 0.10
Standard deviation of our simulation method  (×$(\times 10^{-3})$
@@
Bias of various analytic approximation methods  (×$(\times 10^{-3})$
@@
Bias of our simulation method   (×$(\times 10^{-3})$
$h=1$0.14 -0.30 -0.42 -0.43 -0.43 -0.40 -0.30
@@
Stdev of our simulation method   (×$(\times 10^{-3})$
@@
Bias of various analytic approximation methods   (×$(\times 10^{-3})$
@@ 5.2 Accuracy and speed trade-off
In this subsection, we demonstrate the computational efficiency of our method by comparing the accuracy-speed trade-off with the earlier simulation methods, such as the Euler, low-bias (Chen et al., 2012), and piecewise semi-exact (PSE) (Cai et al., 2017) schemes. For the purpose, we use Cases III and IV. These cases have been tested by Cai et al. (2017), and we take advantage of the extensive records of the CPU time reported therein.
@@
Error (×$(\times 10^{-3})$ and CPU time of our simulation method
@@
Bias (×$(\times 10^{-3})$ and CPU time of the Euler scheme reported in Cai et al. (2017)
$h=1/400$ 1.6 1.5 1.5 1.4 1.3 1.2 49.4
$h=1/800$ 0.7 0.6 0.5 0.5 0.4 0.3 99.1
$h=1/1600$0.3 -0.3 -0.3 -0.3 -0.3 -0.3 194.0
Bias (×$(\times 10^{-3})$ and CPU time of the low-bias scheme reported in Cai et al. (2017)
$h=1/4$ 0.5 0.5 0.5 0.4 0.4 0.4 78.4
$h=1/8$ 0.4 0.4 0.4 0.3 0.3 0.2 175.8
Bias (×$(\times 10^{-3})$ and CPU time of the PSE scheme reported in Cai et al. (2017)
$h=1$ 0.1 0.2 0.2 0.2 0.2 0.2 98.3
@@
The results for Case III in Table 4 show that our method is at least several hundred times faster than the competitors while the bias is much lower. Note that the correlation is zero in this case. When $\rho=0$, both our CEV approximation and Islah’s approximation become exact, and the source of bias is the accuracy of sampling $I_{t}^{h}$. The near-zero error of our method validates our SLN approximation for $I_{t}^{h}$. For the Euler and low-bias schemes, it is necessary to decrease the time step $h$ for more accurate sampling of $I_{t}^{h}$, resulting with slower computation. Although the PSE scheme exhibits good accuracy for single step ($h=1$), it requires prohibitive computation for numerical inverse transform for sampling $I_{t}^{h}$.
$N$ $h$ RMS error (×$\times 10^{-3}$) Time (s)
@@
Table 5: The root mean square (RMS) error and CPU time from Case IV when the number of paths $N$ is doubled and time step $h$ is halved.
@@
RMS $\text{RMS error}=\sqrt{\text{Stdev}^{2}+\text{Bias}^{2}}.$ (21)
@@ 5.3 Comparison with Islah’s approximation
In this last numerical experiment using Case V, we demonstrate superiority of our CEV approximation of conditional $F_{T}$ in Section 4.1 to Islah’s approximation which has been commonly adopted in other schemes. The parameter set is also taken from Cai et al. (2017, Figure 4). We price the at-the-money European option for maturity, $T=1,2,\ldots,10$ with time step $h=1/2$ or 1. On one hand, our simulation scheme consists of (i) SLN sampling of $I_{t}^{h}$ and (ii) CEV model sampling of conditional $F_{T}$. On the other hand, we replace the second step with Islah’s approximation (Proposition 8).
@@
Figure 3 compares the results from the two approaches: our conditional distribution of $F_{T}$ and Islah’s approximation. Figure 3(a) shows the error of the terminal price $E(F_{T})$, which should be equal to $F_{0}=1.1$ in theory. As expected, the results show that our approach preserves the martingale property fairly well for all time-to-maturity $T$. That is, $E(F_{T})=F_{0}$ holds with very high accuracy regardless of $h$. In Islah’s approach, however, $E(F_{T})$ deviates from $F_{0}$. In particular, the deviation accumulates as the time-to-maturity $T$ becomes longer. The time step $h$ has to be smaller to reduce the deviation.
@@ 6 Conclusion
The stochastic-alpha-beta-rho (SABR) model proposed by Hagan et al. (2002) has been widely adopted for pricing derivatives under the stochastic volatility framework. The corresponding simulation methods have been studied extensively. The naive time-discretization schemes, such as the Euler or Milstein schemes, suffer from non-negligible bias caused by the truncation of negative forward values even when a small time step is used. Although researchers proposed various simulation schemes that overcome the limitation of the time-discretization schemes, the schemes have their own limitations. In sampling conditional average variance, the schemes are riddled with either heavy computation (Cai et al., 2017; Leitao et al., 2017a) or inaccurate approximation over a large simulation step (Chen et al., 2012). In sampling conditional forward price, almost all methods adopt Islah (2009)’s NCX2 approximation that neither preserves martingale property nor provides efficient sampling algorithm.
@@
Funding
The works of Lilian Hu and Yue Kuen Kwok are supported by the Guangzhou-HKUST (GZ) Joint Funding Program (Number 2024A03J0630).
@@ Appendix A Conditional moments of $I_{t}^{h}$
The conditional average variance $I_{t}^{h}(\hat{Z})$ is closely related to the exponential functional of BM, defined by
@@
where $Z_{s}$ is a standard Brownian motion. The derivation of Proposition 2 is based on the analytic conditional moments of the exponential functional of BM (Matsumoto and Yor, 2005, (5.4)):
$E\left(\left(A_{t}\right)^{n}|Z_{t}=x\right)=\frac{e^{nx}}{n!t}\int_{x}^{\infty}be^{(x^{2}-b^{2})/2t}\left[\cosh(b)-\cosh(x)\right]^{n}\mathrm{d}b.$
The conditional average variance $I_{t}^{h}(\hat{Z})$ is expressed by $A_{t}$ conditional on the terminal location of the BM, $Z_{t}$:
@@
Therefore, the conditional moments admit the following integral form:
@@
$=\frac{e^{k\hat{\nu}\hat{Z}}}{k!\,\hat{\nu}^{2k}}\int_{\hat{Z}}^{\infty}se^{(\hat{Z}^{2}-s^{2})/2}\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}\hat{Z})\right]^{k}\mathrm{d}s$
$=\frac{e^{k\hat{\nu}\hat{Z}}}{(k-1)!\,\hat{\nu}^{2k-1}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}\hat{Z})\right]^{k-1}\mathrm{d}s,$
@@
$\int_{z}^{\infty}e^{-s^{2}/2}\sinh(as)\,\mathrm{d}s=\frac{N(-z+a)-N(-z-a)}{2n(a)}=\frac{N(z+a)-N(z-a)}{2n(a)},$
we obtain for $k\neq 0$,
$\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(k\hat{\nu}s)\,\mathrm{d}s=k\hat{\nu}\,m_{k}(\hat{Z})\quad\text{where}\quad m_{k}(\hat{Z}):=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
which will be useful in the derivation below. Here, we define $m_{k}(\hat{Z})$ in such a way that $m_{k}(\hat{Z})\to 1$ for any non-zero $k$ as $\hat{\nu}\to 0$.
Next, we evaluate $E((I_{t}^{h})^{k})$. For $k=1$, we obtain
$\mu=\frac{e^{\hat{\nu}\hat{Z}}}{\hat{\nu}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\,\mathrm{d}s=e^{\hat{\nu}\hat{Z}}\,m_{1}(\hat{Z})=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)\,m_{1}.$
For $k=2$, using $\sinh(x)\cosh(x)=\frac{1}{2}\sinh(2x)$, we have
$\mu^{\prime}_{2}$ $=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]\,\mathrm{d}s$
$=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left[\frac{\sinh(2\hat{\nu}s)}{2}-\sinh(\hat{\nu}s)\cosh(\hat{\nu}z)\right]\,\mathrm{d}s$
$=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left[m_{2}-cm_{1}\right].$
For $k=3$, using $\sinh(x)\cosh^{2}(x)=\frac{1}{4}\left[\sinh(3x)+\sinh(x)\right]$, we obtain
$\mu^{\prime}_{3}$ $=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]^{2}\,\mathrm{d}s$
$=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left[\frac{\sinh(3\hat{\nu}s)+\sinh(\hat{\nu}s)}{4}-\sinh(2\hat{\nu}s)\cosh(\hat{\nu}z)+\sinh(\hat{\nu}s)\cosh^{2}(\hat{\nu}z)\right]\,\mathrm{d}s$
@@
For $k=4$, using $\sinh(x)\cosh^{3}(x)=\frac{1}{8}\left[\sinh(4x)+2\sinh(2x)\right]$, we have
$\mu^{\prime}_{4}$ $=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]^{3}\,\mathrm{d}s$
$=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left\{\frac{\sinh(4\hat{\nu}s)+2\sinh(2\hat{\nu}s)}{8}-\frac{3}{4}\left[\sinh(3\hat{\nu}s)+\sinh(\hat{\nu}s)\right]\cosh(\hat{\nu}z)\right.$
$\hskip 120.00018pt\left.+\,\frac{3}{2}\sinh(2\hat{\nu}s)\cosh^{2}(\hat{\nu}z)-\sinh(\hat{\nu}s)\cosh^{3}(\hat{\nu}z)\right\}\,\mathrm{d}s$
@@
$v=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3}),\quad s=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}),\quad\text{and}\quad\kappa=\frac{276}{35}\hat{\nu}^{2}+O(\hat{\nu}^{4}).$

### v2

@@ 5.2 Comparison to analytic approximations
Table 3 shows the five parameter sets used in our numerical tests. In this subsection, we test the accuracy of our algorithm (Algorithm 4) using Cases I and II in Table 3. Tables 4 and 5 show the numerical results for the two sets. The two examples are previously tested in Antonov and Spector (2012) to demonstrate the performance of the cutting-edge analytic approximations for European vanilla options, map to the zero correlation (ZC Map, and hybrid map to the zero correlation (Hyb ZC Map), against the widely used Hagan et al. (2002)’s implied volatility formula. [^3] Since these approximation formulas are all expressed in terms of implied volatilities, we use the Black-Scholes formula to obtain the call option price. Antonov and Spector (2012) use MC method for the benchmark option price. Instead, we use FDM price for the benchmark, which is more precise. Therefore, we also display those option prices for comparison.
Tables 4 and 5 demonstrate that our simulation scheme is highly accurate. As we decrease the time step $h$ from 1 to $1/16$, the option price converges to its true price. Even for a large time step ($h=1$), the option prices are accurate almost up to three decimal points. Such bias level is comparable to the advanced analytic methods. Overall speaking, we note that the bias in Table 5 (Case II) is lower than that in Table 4 (Case I). This is probably due to the observation that the geometric BM approximation for the conditional mean $\bar{F}_{0}^{T}$ in Eq. (13) holds better as $\beta=0.6$ (Case II) is closer to $\beta=1$ (lognormal SABR model) than $\beta=0.3$ (Case I).
@@
Bias of our simulation method  $(\times 10^{-3})$
$h=1$ -1.22 -1.49 -0.37 0.49 1.28 1.72 1.32
$h=1/4$ -0.46 -0.24 0.22 0.42 0.56 0.56 0.48
$h=1/16$ -0.34 -0.20 0.00 0.05 0.11 0.10 0.10
Standard deviation of our simulation method  $(\times 10^{-3})$
@@
Bias of various analytic approximation methods  $(\times 10^{-3})$
@@
Bias of our simulation method   $(\times 10^{-3})$
$h=1$ -0.14 -0.30 -0.42 -0.43 -0.43 -0.40 -0.30
@@
Stdev of our simulation method   $(\times 10^{-3})$
@@
Bias of various analytic approximation methods   $(\times 10^{-3})$
@@ 5.3 Accuracy and speed trade-off
In this subsection, we demonstrate the computational efficiency of our method by comparing the accuracy-speed trade-off with the earlier simulation methods, such as the Euler, low-bias (Chen et al., 2012), and piecewise semi-exact (PSE) (Cai et al., 2017) schemes. We use Cases III and IV to achieve better comparison since these cases have been tested by Cai et al. (2017). We take advantage of the extensive records of the CPU time reported therein.
@@
Error $(\times 10^{-3})$ and CPU time of our simulation method
@@
Bias $(\times 10^{-3})$ and CPU time of the Euler scheme reported in Cai et al. (2017)
$h=1/400$ 1.6 1.5 1.5 1.4 1.3 1.2 4.12
$h=1/800$ 0.7 0.6 0.5 0.5 0.4 0.3 8.26
$h=1/1600$ -0.3 -0.3 -0.3 -0.3 -0.3 -0.3 16.2
Bias $(\times 10^{-3})$ and CPU time of the low-bias scheme reported in Cai et al. (2017)
$h=1/4$ 0.5 0.5 0.5 0.4 0.4 0.4 6.54
$h=1/8$ 0.4 0.4 0.4 0.3 0.3 0.2 14.7
Bias $(\times 10^{-3})$ and CPU time of the PSE scheme reported in Cai et al. (2017)
$h=1$ 0.1 0.2 0.2 0.2 0.2 0.2 8.19
@@
The results for Case III in Table 6 show that our method is at least one hundred times faster than the competitors while the bias is much lower. Under $\rho=0$ in this case, both our CEV approximation and Islah’s approximation become exact, and the source of bias is the accuracy of sampling $I_{t}^{h}$. The near-zero error of our method validates extremely high accuracy of our SLN approximation for $I_{t}^{h}$. For the Euler and low-bias schemes, it is necessary to decrease the time step $h$ for more accurate sampling of $I_{t}^{h}$, resulting with slower computation. Although the PSE scheme exhibits good accuracy for single step ($h=1$), it requires prohibitive computation for numerical inverse transform for sampling $I_{t}^{h}$.
NN $h$ RMS error ($\times 10^{-3}$) Time (s)
@@
Table 7: The root mean square (RMS) error and CPU time from Case IV when the number of paths NN is doubled and time step $h$ is halved.
@@
$\text{RMS error}=\sqrt{\text{Stdev}^{2}+\text{Bias}^{2}}.$ (26)
@@ 5.4 Comparison with Islah’s approximation
In this last numerical experiment using Case V, we demonstrate superiority of our CEV approximation of conditional $F_{T}$ in Section 4.1 to Islah’s approximation which has been commonly adopted in other simulation schemes. The parameter set is also taken from Cai et al. (2017, Figure 4). We price the at-the-money European option for maturity, $T=1,2,\ldots,10$ with time step $h=1/2$ or 1. On one hand, our simulation scheme consists of (i) SLN sampling of $I_{t}^{h}$ and (ii) CEV model sampling of conditional $F_{T}$. On the other hand, we replace the second step with Islah’s approximation (Eq. (25)).
@@
Figure 3 compares the results from the two approaches: our conditional distribution of $F_{T}$ and Islah’s approximation. Figure 3(a) shows the error of the terminal price $E(F_{T})$, which should be equal to $F_{0}=1.1$ in theory. As expected, the results show that our approach preserves the martingale property of the forward price fairly well for all time-to-maturity $T$. That is, $E(F_{T})=F_{0}$ holds with very high accuracy regardless of $h$. In Islah’s approach, however, $E(F_{T})$ deviates from $F_{0}$. In particular, the deviation from $F_{0}$ accumulates as the time-to-maturity $T$ becomes longer. The time step $h$ has to be smaller to reduce the deviation.
@@ 6 Conclusion
The stochastic-alpha-beta-rho (SABR) model proposed by Hagan et al. (2002) has been widely adopted for pricing derivatives under the stochastic volatility framework. The corresponding simulation methods have been studied extensively. The naive time-discretization schemes, such as the Euler or Milstein schemes, suffer from non-negligible bias caused by the truncation of negative forward values even when a small time step is used. Although researchers proposed various simulation schemes that overcome the limitation of the time-discretization schemes, the schemes have their own limitations. In sampling conditional average variance, the simulation schemes are riddled with either heavy computation (Cai et al., 2017; Leitao et al., 2017a) or inaccurate approximation over a large simulation step (Chen et al., 2012). In sampling conditional forward price, almost all simulation algorithms adopt Islah (2009)’s NCX2 approximation that neither preserves martingale property nor provides efficient sampling algorithm.
@@
Acknowledgment
The works of Lilian Hu and Yue Kuen Kwok are supported by the Guangzhou-HKUST (GZ) Joint Funding Program (Number 2024A03J0630). The work of Yue Kuen Kwok is also supported by Project 2023CX10X1.
@@ Appendix A Conditional moments of $I_{t}^{h}$
[…]
@@
where $Z_{s}$ is a standard BM. The derivation of Proposition 2 is based on the analytic conditional moments of the exponential functional of BM (Matsumoto and Yor, 2005, (5.4)):
$E\left(\left(A_{t}\right)^{n}|Z_{t}=x\right)=\frac{e^{nx}}{n!t}\int_{x}^{\infty}be^{(x^{2}-b^{2})/2t}\left(\cosh b-\cosh x\right)^{n}\mathrm{d}b.$
[…]
@@
Accordingly, the conditional moments of $I_{t}^{h}$ admit the following integral form:
@@
$=\frac{e^{k\hat{\nu}\hat{Z}}}{k!\,\hat{\nu}^{2k}}\int_{\hat{Z}}^{\infty}se^{(\hat{Z}^{2}-s^{2})/2}\left(\cosh\hat{\nu}s-\cosh\hat{\nu}\hat{Z}\right)^{k}\mathrm{d}s$
$=\frac{e^{k\hat{\nu}\hat{Z}}}{(k-1)!\,\hat{\nu}^{2k-1}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\left(\cosh\hat{\nu}s-\cosh\hat{\nu}\hat{Z}\right)^{k-1}\mathrm{d}s,$
@@
$\int_{z}^{\infty}e^{-s^{2}/2}\sinh as\,\mathrm{d}s=\frac{N(-z+a)-N(-z-a)}{2n(a)}=\frac{N(z+a)-N(z-a)}{2n(a)},$
for $k\neq 0$, we obtain
$\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh k\hat{\nu}s\,\mathrm{d}s=k\hat{\nu}\,m_{k}(\hat{Z})\quad\text{where}\quad m_{k}(\hat{Z}):=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
[…]
Next, we evaluate the moments $E((I_{t}^{h})^{k})$, $k=1,2,3,4$. For $k=1$, we obtain
$\mu=\frac{e^{\hat{\nu}\hat{Z}}}{\hat{\nu}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\,\mathrm{d}s=e^{\hat{\nu}\hat{Z}}\,m_{1}(\hat{Z})=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)\,m_{1}.$
For $k=2$, using $\sinh x\cosh x=\frac{1}{2}\sinh 2x$, we have
$\mu^{\prime}_{2}$ $=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\left(\cosh\hat{\nu}s-\cosh\hat{\nu}z\right)\,\mathrm{d}s$
$=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left(\frac{\sinh 2\hat{\nu}s}{2}-\sinh\hat{\nu}s\cosh\hat{\nu}z\right)\,\mathrm{d}s$
$=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left(m_{2}-cm_{1}\right).$
For $k=3$, using $\sinh x\cosh^{2}x=\frac{1}{4}\left(\sinh 3x+\sinh x\right)$, we obtain
$\mu^{\prime}_{3}$ $=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\left(\cosh\hat{\nu}s-\cosh\hat{\nu}z\right)^{2}\,\mathrm{d}s$
$=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left(\frac{\sinh 3\hat{\nu}s+\sinh\hat{\nu}s}{4}-\sinh 2\hat{\nu}s\cosh\hat{\nu}z+\sinh\hat{\nu}s\cosh^{2}\hat{\nu}z\right)\,\mathrm{d}s$
@@
For $k=4$, using $\sinh x\cosh^{3}x=\frac{1}{8}\left(\sinh 4x+2\sinh 2x\right)$, we have
$\mu^{\prime}_{4}$ $=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\left(\cosh\hat{\nu}s-\cosh\hat{\nu}z\right)^{3}\,\mathrm{d}s$
$=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left[\frac{\sinh 4\hat{\nu}s+2\sinh 2\hat{\nu}s}{8}-\frac{3}{4}\left(\sinh 3\hat{\nu}s+\sinh\hat{\nu}s\right)\cosh\hat{\nu}z\right$.
$\hskip 120.00018pt\left.+\,\frac{3}{2}\sinh 2\hat{\nu}s\cosh^{2}\hat{\nu}z-\sinh\hat{\nu}s\cosh^{3}\hat{\nu}z\right]\,\mathrm{d}s$
@@
$Cv=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3}),\quad s=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}),\quad\text{and}\quad\kappa=\frac{276}{35}\hat{\nu}^{2}+O(\hat{\nu}^{4}).$